### v1

@@ 5.1 Logic Strategy Maps
The "Strategy Maps" in Figure 6 visualize the output of the gating network $G(E)$. In homogeneous regions (e.g., the sky in Figure 5(b)), the gate values approach 0, activating the Context Expert which uses dilated convolutions to integrate broad spatial information and suppress noise. Conversely, at object boundaries (e.g., the bear’s outline), gate values approach 1, activating the Boundary Expert to preserve sharp pixel transitions via pointwise convolutions. This confirms the network actively switches strategies based on local image topology, rather than applying a "one-size-fits-all" operation.
@@ 5.2 Fuzzy Rule Firing Maps
The "Rule Firing Maps" reveal which logical premises are active. By inspecting the learned centers and consequents, we can semantically label the rules:
  • Rule 1 (Strong Edge): Centers at high $x_{1}$ and $x_{2}$. Fires when both Sobel gradient and U-Net confidence are high, representing "Unambiguous Edges."
  • Rule 2 (Semantic Edge): Center at low $x_{1}$ but high $x_{2}$. Fires when gradients are weak but the U-Net is confident, capturing "Soft Boundaries" (e.g., low-contrast contours).
  • Rule 3 (Noise Suppression): Centers at low values. Detects background and actively suppresses it.
  • Rule 4 (Texture Suppression): Centers at high $x_{1}$ but low $x_{2}$. Crucially, the consequent weight is negative (Fig. 8(c)), explicitly suppressing false positive gradients to clean up the output.
While TSK systems are universal approximators capable of fitting functions via piecewise segments, the explainability here arises from sparsity and semantic grounding. Unlike naive approximations requiring thousands of segments, we condense the U-Net’s decision boundary into exactly four rules. Furthermore, by constraining inputs to interpretable axes (Gradient $x_{1}$, Semantic Confidence $x_{2}$), the learned weights reveal the specific mechanistic logic used to resolve conflicts, rather than simply providing a black-box mathematical fit.

### v2

@@ 5.1 Logic Strategy Maps
The "Strategy Maps" in Figure 6 visualize the output of the gating network $G(E)$.
  • Observation: In homogeneous regions (e.g., the sky in Fig 5), the gate values approach 0, activating the Context Expert. This expert uses dilated convolutions to integrate broad spatial information, effectively suppressing noise.
  • Observation: At object boundaries (e.g., the bear’s outline), the gate values approach 1, activating the Boundary Expert. This expert uses pointwise convolutions to preserve sharp pixel transitions.
  • Significance: This proves the network is actively switching strategies based on local image topology, rather than applying a "one-size-fits-all" convolution.
@@ 5.2 Fuzzy Rule Firing Maps
The "Rule Firing Maps" reveal which logical premises are active. By inspecting the learned centers and consequents in the figure, we can semantically label the rules:
  • Rule 1 (Strong Edge): Centers at high $x_{1}$ and $x_{2}$. This rule fires when both Sobel gradient and U-Net confidence are high. It represents "Unambiguous Edges."
  • Rule 2 (Semantic Edge): Center at low $x_{1}$ but high $x_{2}$. This rule fires when the gradient is weak, but the U-Net is confident. This captures "Soft Boundaries" (e.g., visual contours with low contrast).
  • Rule 3 (Noise Suppression): Centers at low values. This rule detects background and actively suppresses it.
  • Rule 4 (Texture Suppression) : Centers at high $x_{1}$ (Edge Strength) but low $x_{2}$ (Semantic Confidence). The consequent weight for this rule is negative (as seen in Figure 8(c)), effectively suppressing false positive gradients to clean up the final output. For a deeper quantitative analysis of how these rules are formed, refer to the trained rules visualization in Figure 8(a) and 8(b).
This level of transparency allows domain experts to verify why an edge was detected. For example, if an edge is detected via Rule 2, the user knows it relies on learned semantic context rather than raw pixel intensity.